OpenStudy (anonymous):

On Monday, the office staff at your school paid $8.77 for 4 cups of coffee and 7 bagels. On Wednesday, they paid $15.80 for 8 cups of coffee and 14 bagels. Can you determine the cost of a bagel? Explain.

OpenStudy (anonymous):

that would be 18 right

OpenStudy (anonymous):

Could you explain ?

OpenStudy (anonymous):

You said: On Monday, the office staff at your school paid $8.77 for 4 cups of coffee and 7 bagels. On Wednesday, they paid $15.80 for 8 cups of coffee and 14 bagels. Can you determine the cost of a bagel? Explain. so that means the cost of a beagle is $18

OpenStudy (anonymous):

The bagel would be less than the total cost.

hero (hero):

4c + 7b = 8.77 8c + 14b = 15.80 Solve for b

OpenStudy (anonymous):

Thanks . I got it .

hero (hero):

@isaidrawr , I got your back :D

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

4c + 7b = 8.77 8c + 14b = 15.80

hero (hero):

@Arieboo11 , I posted that already